# Self-Organizing Dual-Buffer Adaptive Clustering Experience Replay (SODASER) for Safe Reinforcement Learning in Optimal Control

### v1

@@ Abstract
This paper proposes a novel reinforcement learning framework, named Self-Organizing Dual-buffer Adaptive Clustering Experience Replay (SODACER), designed to achieve safe and scalable optimal control of nonlinear systems. The proposed SODACER mechanism consisting of a Fast-Buffer for rapid adaptation to recent experiences and a Slow-Buffer equipped with a self-organizing adaptive clustering mechanism to maintain diverse and non-redundant historical experiences. The adaptive clustering mechanism dynamically prunes redundant samples, optimizing memory efficiency while retaining critical environmental patterns. The approach integrates SODASER with Control Barrier Functions (CBFs) to guarantee safety by enforcing state and input constraints throughout the learning process. To enhance convergence and stability, the framework is combined with the Sophia optimizer, enabling adaptive second-order gradient updates. The proposed SODACER-Sophia’s architecture ensures reliable, effective, and robust learning in dynamic, safety-critical environments, offering a generalizable solution for applications in robotics, healthcare, and large-scale system optimization. The proposed approach is validated on a nonlinear Human Papillomavirus (HPV) transmission model with multiple control inputs and safety constraints. Comparative evaluations against random and clustering-based experience replay methods demonstrate that SODACER achieves faster convergence, improved sample efficiency, and a superior bias-variance trade-off, while maintaining safe system trajectories, validated via the Friedman test.
@@ Introduction
While RL has shown success across various domains, achieving safe, optimal control for nonlinear systems with state and input constraints remains challenging. Techniques such as Lyapunov-based methods and constraint-satisfaction criteria have been incorporated into RL to address safety concerns [5, 6, 7]. However, these approaches face limitations when applied to high-dimensional systems with rapidly evolving dynamics. Furthermore, they often struggle with the bias-variance trade-off inherent in RL, where prioritizing stability may reduce adaptability. Experience Replay (ER) has become instrumental in addressing these RL challenges, enhancing learning efficiency and adaptability by retaining and reusing past interactions [8].
Experience replay fundamentally improves RL efficiency by storing past experiences, which reduces the need for continuous system interaction a significant benefit for applications where real-time experimentation is impractical [9, 10]. While traditional ER methods utilize random sampling, these approaches can fall short in nonstationary environments or complex control tasks. Schaul et al. [11] proposed Prioritized Experience Replay (PER), which enhances sample efficiency by assigning higher sampling probabilities to experiences with larger temporal-difference errors. Although PER improves focus on crucial transitions, it risks overemphasizing outliers or outdated samples, potentially leading to instability in changing environments.
@@
To optimize computational efficiency within this framework, we employ the Sophia optimizer [23], known for its dynamic step adjustments that improve convergence rates. For systems experiencing frequent or severe disturbances, we advise adopting a robust, dynamic optimization solution like ET-DSIWO [24]. This integration enhances the adaptability of the SODACER framework in high-dimensional RL applications, further contributing to efficient gradient-based optimization. By combining SODACER with the Sophia optimizer and validating it on a real-world public health model, this study contributes a scalable, robust RL approach for complex, dynamic systems, offering new pathways for effective control in high-dimensional, constraint-driven applications.
We validate our methodology by applying SODACER to the optimal control of a compartmental model for Human Papillomavirus (HPV) transmission dynamics. Given the complex, nonlinear nature of HPV transmission, with inherent state and input constraints, this model serves as a relevant test case for adaptive control policies in a constrained environment. We employed the model and system parameters that had been established previously in our earlier work. Our preceding publication has comprehensively examined and detailed this disease model and convergence study [25]. Our approach addresses the unmet need for RL-based public health strategies, aiming to minimize infection rates and intervention costs through safe, data-driven policies. Prior studies by Saldaña et al. [18] and Malik et al. [19, 20] have underscored the challenges of controlling HPV spread, while Brown et al. [21] highlighted the cost-effectiveness of vaccination programs. This work builds on these insights, demonstrating the potential of safe RL methods for optimizing health policy.
@@
$\mathcal{J}(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}\mathcal{L}(\tau)d\tau,$ (3)
where $Q\in\mathbb{R}^{p\times p}$ is a positive definite matrix, $\nu$ is a discount factor, and $\mathcal{L}(\tau)=e^{\top}(\tau)Qe(\tau)+U(u(\tau))$. The term $U(u)$ represents the cost associated with control efforts and is given by:
$U(u)=2\sum_{i=1}^{q}\int_{0}^{u_{i}}\left(\kappa_{i}\Phi_{i}\tanh^{-1}\left(\frac{\theta_{i}}{\kappa_{i}}\right)\right)^{T}d\theta_{i},$ (4)
where $\Phi_{i}>0$ for all $i=1,\dots,q$.
@@
$\mathcal{H}(e,u,x,\mathcal{J}_{e})=e^{\top}(t)Qe(t)+U(u)-\nu\mathcal{J}(e)$
$+\mathcal{J}_{e}^{\top}\left(f(x(t))+g(x(t))u(t)\right),$ (5)
where $\mathcal{J}_{e}\triangleq\frac{\partial\mathcal{J}(e)}{\partial e}$ represents the gradient of the objective function with respect to the error $e(t)$.
By applying the stationarity condition $\frac{\partial\mathcal{H}(e,u,x,\mathcal{J}_{e})}{\partial u}=0$, the optimal control input $u^{*}=[u_{1}^{*},u_{2}^{*},\dots,u_{q}^{*}]^{T}$ is derived as:
$u_{i}^{*}=-\kappa_{i}\tanh\left(\frac{1}{2\kappa_{i}\Phi_{i}}g^{\top}(x)\mathcal{J}_{e}^{*}(x)\right).$ (6)
@@
$e^{\top}(t)Qe(t)+U(u^{*})-\nu\mathcal{J}^{*}(e)$
$+\mathcal{J}_{e}^{*\top}\left(f(x(t))+g(x(t))u^{*}(t)\right)=0.$ (7)
@@
proof The proof follows the methodology outlined in [5].
@@
In addition to optimizing the cost functional, the control $u(t)$ must satisfy state constraints that ensure the system operates within safe boundaries. To ensure the system adheres to the state constraints in a dynamically responsive manner, we employ the concept of Safe Optimal Control through Control Barrier Functions (CBFs). A Control Barrier Function $h(x)$ provides a measure of “safety” by enforcing constraints on the state trajectory $x(t)$ under the influence of the control input $u(t)$. These constraints are represented as:
@@
where $h:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth function defining the permissible state space. The inequality $h(x(t))\geq 0$ ensures that the state remains within a safe, admissible set throughout the control horizon.
@@
Minimize $\mathcal{J}(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}\mathcal{L}(\tau)d\tau,$ (9a)
@@
This inequality implies that, as long as $h(x(0))\geq 0$ at the initial time, the system state $x(t)$ will remain in the safe set $\{x\in\mathbb{R}^{n}:h(x)\geq 0\}$ for all $t\geq 0$.
@@
Minimize $\mathcal{J}(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}\mathcal{L}(\tau)d\tau,$ (11a)
@@
The safe optimal control policy $u^{*}(t)$ is derived by solving the above constrained optimization problem, balancing the objectives of minimizing the cost functional $J$ while satisfying the CBF-based safety constraint. By integrating CBFs into the optimization, the resulting policy inherently ensures safety, enabling the system to achieve performance goals without violating the state constraints.
@@
In order to approximate the objective function $\mathcal{J}(x(t))$ over the compact set $\Omega$, a single-layer adaptive critic neural network (ANN) is employed. This ANN is trained to approximate the value of the objective function as follows:
$\mathcal{J}^{*}(e(t),W^{*})=\varphi^{\top}(e(t))W^{*}+\tilde{\epsilon}(e(t)),$ (12)
where $W(t)\in\mathbb{R}^{l}$ represents the adjustable weight vector of the ANN, $\varphi(e(t))\in\mathbb{R}^{l}$ is the activation function, and $\tilde{\epsilon}(e(t))$ denotes the approximation error of the ANN. The parameter $l$ corresponds to the number of neurons in the hidden layer.
Accordingly, the gradient of $\mathcal{J}^{*}(e(t))$, denoted as $\triangledown_{e}\mathcal{J}$, is expressed as:
$\triangledown_{e}\mathcal{J}^{*}=\triangledown_{e}\varphi^{\top}W+\triangledown_{e}\tilde{\epsilon}$, (13)
@@
Since the error $\tilde{\epsilon}(e(t))$ is not directly measurable, we aim to estimate the weight vector $W$. Hence, the approximations for $\mathcal{J}(e(t))$ and its gradient are rewritten as:
$\hat{\mathcal{J}}(e(t))=\varphi(e(t))^{\top}\hat{W}$, (14)
$\triangledown_{e}\hat{\mathcal{J}}=\triangledown_{e}\varphi^{\top}\hat{W}$, (15)
@@
$\mathcal{J^{*}}-\hat{\mathcal{J}}=\varphi^{\top}(t)(W^{*}-\hat{W})+\tilde{\epsilon}(e(t)),$ (16)
@@
$\mathcal{J^{*}}=\hat{\mathcal{J}}+\tilde{\delta}+\tilde{\epsilon}(e(t)),$ (17)
@@
$\min\hskip 2.84544ptE=\frac{1}{2}\hat{\mathcal{H}}^{T}\hat{\mathcal{H}}$, (20)
Here, the objective function presented in (20) is designed to minimize the quadratic Hamiltonian. By minimizing the objective function $E$, we achieve optimized system performance, characterized by reduced control input $\hat{u}$ while ensuring that the weights $\hat{W}$ converge to their optimal values.
@@
The Fast Buffer is a small, adaptive component that prioritizes recent experiences aligned with the current policy $\pi_{\text{curr}}$. New system interactions generate samples, denoted as $S_{\text{new}}=(\mathbf{x}_{t},\mathbf{u}_{t})$, which are immediately stored in the Fast Buffer. By focusing on data that reflects recent agent-environment interactions, the Fast Buffer provides low-bias, high-variance samples, allowing rapid adaptation to evolving dynamics and aligning the agent’s behavior with current environmental conditions.
@@
where $\mathbf{c}_{j}$ denotes the centroid, and $\sigma_{j}^{2}$ represents the variance of cluster $\mathcal{C}_{j}$. This metric enables efficient clustering by ensuring each sample’s integration with the most representative cluster, preventing over-segmentation and enhancing the coherence of clustered data points. The formation of clusters is guided by the computed membership strengths. For any sample with a low similarity to all existing clusters (i.e., $\max(\mu_{\mathcal{C}}(S_{\text{old}}))\leq\Gamma_{th}$), a new cluster $\mathcal{C}_{\text{new}}$ is created in the Cluster Buffer. This new cluster initializes with $\mathbf{c}_{\text{new}}=S_{\text{old}}$ and a variance $\sigma_{\text{new}}=\sigma_{0}=0.02$. Conversely, if a sample fits well within an existing cluster $\mathcal{C}_{j}$, the cluster center $\mathbf{c}_{j}$ is updated to incorporate the new data point:
$\mathbf{c}_{j}=\frac{\mathcal{N}_{j}\mathbf{c}_{j}+S_{\text{old}}}{\mathcal{N}_{j}+1},$ (22)
@@
$\|\mathbf{c}_{i}-\mathbf{c}_{j}\|<\gamma\max(\sigma_{i},\sigma_{j}),$ (25)
Now if (25) is satisfied, then overlap between dusters $i$ and $j$ is accured and without loosing the generality, let’s assume $\sigma_{i}\geqslant\sigma_{j}$ then the cluster $\mathcal{C}_{j}$ will be pruned and they are merged to form a consolidated cluster $\mathcal{C}_{\text{merged}}$ with its center recalculated as:
$\mathbf{c}_{\text{merged}}=\frac{\mathcal{N}_{i}\mathbf{c}_{i}+\mathcal{N}_{j}\mathbf{c}_{j}}{\mathcal{N}_{i}+\mathcal{N}_{j}}.$ (26)
@@
Remark 1 The value of the parameter $\gamma$, which quantifies the overlap between two clusters, is derived as follows. The overlap is determined by the membership strength $\mu_{\mathcal{C}_{i}}(c_{j})$ between the centroids $c_{i}$ and $c_{j}$ based on (21). To ensure a sufficient degree of overlap, we impose the following condition,
$\text{max}(\mu_{\mathcal{C}_{i}}(c_{i}),\mu_{\mathcal{C}_{j}}(c_{j})))\geq\bar{\gamma}=0.95\%.$
@@
$\frac{\|c_{i}-c_{j}\|^{2}}{2\sigma_{i}^{2}}\geq\ln(\bar{\gamma})\quad\Rightarrow\quad\frac{\|c_{i}-c_{j}\|^{2}}{\sigma_{i}^{2}}\leq-2\ln(\bar{\gamma}).$
Thus, the cluster $\mathcal{C}_{j}$ is covered completely with the cluster $\mathcal{C}_{i}$ if:
$\|c_{i}-c_{j}\|\leq\sqrt{-2\ln(\bar{\gamma})}\cdot\sigma_{i}.$

### v2

# Self-Organizing Dual-Buffer Adaptive Clustering Experience Replay (SODACER) for Safe Reinforcement Learning in Optimal Control

@@ Abstract
This paper proposes a novel reinforcement learning framework, named Self-Organizing Dual-buffer Adaptive Clustering Experience Replay (SODACER), designed to achieve safe and scalable optimal control of nonlinear systems. The proposed SODACER mechanism consisting of a Fast-Buffer for rapid adaptation to recent experiences and a Slow-Buffer equipped with a self-organizing adaptive clustering mechanism to maintain diverse and non-redundant historical experiences. The adaptive clustering mechanism dynamically prunes redundant samples, optimizing memory efficiency while retaining critical environmental patterns. The approach integrates SODACER with Control Barrier Functions (CBFs) to guarantee safety by enforcing state and input constraints throughout the learning process. To enhance convergence and stability, the framework is combined with the Sophia optimizer, enabling adaptive second-order gradient updates. The proposed SODACER-Sophia’s architecture ensures reliable, effective, and robust learning in dynamic, safety-critical environments, offering a generalizable solution for applications in robotics, healthcare, and large-scale system optimization. The proposed approach is validated on a nonlinear Human Papillomavirus (HPV) transmission model with multiple control inputs and safety constraints. Comparative evaluations against random and clustering-based experience replay methods demonstrate that SODACER achieves faster convergence, improved sample efficiency, and a superior bias-variance trade-off, while maintaining safe system trajectories, validated via the Friedman test.
@@ Introduction
While RL has shown success across various domains, achieving safe, optimal control for nonlinear systems with state and input constraints remains challenging. Techniques such as Lyapunov-based methods and constraint-satisfaction criteria have been incorporated into RL to address safety concerns [5, 6, 7]. However, these approaches face limitations when applied to high-dimensional systems with rapidly evolving dynamics. This is mainly due to the curse of dimensionality, increased data nonstationarity, and the difficulty of maintaining an effective balance between stability and adaptability under rapidly changing system behaviors. Furthermore, they often struggle with the bias-variance trade-off inherent in RL, where prioritizing stability may reduce adaptability. Experience Replay (ER) has become instrumental in addressing these RL challenges, enhancing learning efficiency and adaptability by retaining and reusing past interactions [8].
Experience replay fundamentally improves RL efficiency by storing past experiences, which reduces the need for continuous system interaction a significant benefit for applications where real-time experimentation is impractical [9, 10]. While traditional ER methods utilize random sampling, these approaches may fall short in nonstationary environments or complex control tasks. Schaul et al. [11] proposed Prioritized Experience Replay (PER), which enhances sample efficiency by assigning higher sampling probabilities to experiences with larger temporal-difference errors. Although PER improves focus on crucial transitions, it risks overemphasizing outliers or outdated samples, potentially leading to instability in changing environments.
@@
Recent studies have explored multi‑buffer and parallel sampling strategies to enhance the efficiency of off‑policy reinforcement learning. For instance, Chen et al. [18] proposed a double sampling mechanism combined with a mirror experience pool to accelerate training in multi‑agent systems. More recently, attention‑adjusted prioritized experience replay was introduced in [19] to refine sample weighting during optimization. Although these approaches improve data efficiency and learning performance, they rely on static replay structures and primarily operate at the sampling or loss‑adjustment level. In contrast, the proposed SODACER framework introduces a functionally asymmetric dual‑buffer architecture with a self‑organizing adaptive clustering mechanism, explicitly designed to address non‑stationarity and safety‑critical requirements in nonlinear optimal control problems.
@@
To optimize computational efficiency within this framework, we employ the Sophia optimizer [25], known for its dynamic step adjustments that improve convergence rates. For systems experiencing frequent or severe disturbances, we advise adopting a robust, dynamic optimization solution like ET-DSIWO [26]. This integration enhances the adaptability of the SODACER framework in high-dimensional RL applications, further contributing to efficient gradient-based optimization. By combining SODACER with the Sophia optimizer and validating it on a real-world public health model, this study contributes a scalable, robust RL approach for complex, dynamic systems, offering new pathways for effective control in high-dimensional, constraint-driven applications. Rather than reducing the system dimension explicitly, SODACER addresses high‑dimensionality by improving experience utilization through a dual‑buffer structure and self‑organizing clustering, enabling efficient learning in large state spaces.
Recent methodological advances have expanded the toolkit for safe reinforcement learning. For instance, comprehensive surveys of certificate-based methods [27] demonstrate how neural Lyapunov, barrier, and contraction techniques provide formal safety guarantees in robotics and control, forming a theoretical foundation for our use of Control Barrier Functions. Concurrently, continual learning frameworks [28] have been developed to maintain reliable performance under non-stationary conditions, directly addressing the need for adaptive experience management in dynamic environments. Building upon these developments, our work integrates adaptive safety certification with a self-organizing experience replay mechanism to achieve both safety and efficiency in nonlinear optimal control.
We validate our methodology by applying SODACER to the optimal control of a compartmental model for Human Papillomavirus (HPV) transmission dynamics. Given the complex, nonlinear nature of HPV transmission, with inherent state and input constraints, this model serves as a relevant test case for adaptive control policies in a constrained environment. We employed the model and system parameters that had been established previously in our earlier work. Our preceding publication has comprehensively examined and detailed this disease model and convergence study [29]. Our approach addresses the unmet need for RL-based public health strategies, aiming to minimize infection rates and intervention costs through safe, data-driven policies. Prior studies by Saldaña et al. [20] and Malik et al. [21, 22] have underscored the challenges of controlling HPV spread, while Brown et al. [23] highlighted the cost-effectiveness of vaccination programs. This work builds on these insights, demonstrating the potential of safe RL methods for optimizing health policy. In this work, safety is enforced by integrating the proposed SODACER framework with Control Barrier Functions (CBFs), which act as an online safety filter to correct the policy-generated control actions and guarantee forward invariance of the predefined safe set during learning and execution.
@@
$J(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}L(\tau)d\tau,$ (3)
where $Q\in\mathbb{R}^{p\times p}$ is a positive definite matrix, $\nu$ is a discount factor, and $L(\tau)=e^{\top}(\tau)Qe(\tau)+U(u(\tau))$. The term $U(u)$ represents the cost associated with control efforts and is given by:
$U(u)=2\sum_{i=1}^{q}\int_{0}^{u_{i}}\kappa_{i}\Phi_{i}\tanh^{-1}\left(\frac{\theta_{i}}{\kappa_{i}}\right)d\theta_{i},$ (4)
Here, $\Phi_{i}>0$ for all $i=1,\dots,q$ denotes a control weighting parameter associated with the $i$-th control input. This parameter regulates the relative contribution of each control channel to the control effort cost and ensures appropriate scaling in the saturated control structure induced by the hyperbolic tangent function. Larger values of $\Phi_{i}$ impose a higher penalty on the corresponding control input, leading to more conservative control actions, while smaller values allow more aggressive control behavior.
Here, $U(u)$ is chosen as a positive definite control effort penalty to regularize the optimization problem and prevent excessive or unsafe control actions.
@@
$H(e,u,x,J_{e})=e^{\top}(t)Qe(t)+U(u)-\nu J(e)$
$+J_{e}^{\top}\left(f(x(t))+g(x(t))u(t)\right),$ (5)
where $J_{e}\triangleq\frac{\partial J(e)}{\partial e}$ represents the gradient of the objective function with respect to the error $e(t)$.
By applying the stationarity condition $\frac{\partial H(e,u,x,J_{e})}{\partial u}=0$, the optimal control input $u^{*}=[u_{1}^{*},u_{2}^{*},\dots,u_{q}^{*}]^{\top}$ is derived as:
$u_{i}^{*}=-\kappa_{i}\tanh\left(\frac{1}{2\kappa_{i}\Phi_{i}}g^{\top}(x)J_{e}^{*}\right).$ (6)
@@
$e^{\top}(t)Qe(t)+U(u^{*})-\nu J^{*}(e)$
$+J_{e}^{*\top}\left(f(x(t))+g(x(t))u^{*}(t)\right)=0.$ (7)
@@
proof: The proof follows the methodology outlined in [5].
Remark: Although the HJB equation is approximated using neural networks, closed-loop stability is guaranteed in the sense of uniform ultimate boundedness, while safety is ensured independently via control barrier function constraints.
@@
In addition to optimizing the cost functional, the control $u(t)$ must satisfy state constraints that ensure the system operates within safe boundaries. To ensure the system adheres to the state constraints in a dynamically responsive manner, we employ the concept of safe optimal control through CBFs. A Control Barrier Function $h(x)$ provides a measure of “safety” by enforcing constraints on the state trajectory $x(t)$ under the influence of the control input $u(t)$. The inequality $h(x(t))\geq 0$ ensures that the state remains within a safe, admissible set throughout the control horizon. These constraints are represented as:
@@
where $h:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth function defining the permissible state space.
Remark: In the context of neural-network-based approximated dynamic programming, the approximation residual $\tilde{\varepsilon}(e(t))$ is assumed to be small enough, according to the universal approximation theorem.
@@
Minimize $J(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}L(\tau)d\tau,$ (9a)
@@
The modified control problem incorporating the CBF constraint becomes:
Minimize $J(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}L(\tau)d\tau,$ (11a)
@@
The safe optimal control policy $u^{*}(t)$ is derived by solving the above constrained optimization problem, balancing the objectives of minimizing the cost functional $J$ while satisfying the CBF-based safety constraint. By integrating CBFs into the optimization, the resulting policy inherently ensures safety, enabling the system to achieve performance goals without violating the state constraints. In the proposed framework, the reinforcement learning policy generates a nominal control input based on the observed system state. This nominal action is then passed through a CBF-based safety filter, which solves a constrained optimization problem to minimally modify the action when necessary, ensuring that the CBF condition is satisfied. As a result, all executed control inputs preserve the forward invariance of the safe set, while the learning agent remains focused on optimizing performance objectives.
@@
In order to approximate the objective function $J(x(t))$ over the compact set $\Omega$, a single-layer adaptive critic neural network (ANN) is employed. This ANN is trained to approximate the value of the objective function as follows:
$J^{*}(e(t),W^{*})=\varphi^{\top}(e(t))W^{*}+\tilde{\epsilon}(e(t)),$ (12)
where $W(t)\in\mathbb{R}^{l}$ represents the adjustable weight vector of the ANN, $\varphi(e(t))\in\mathbb{R}^{l}$ is the activation function, and $\tilde{\epsilon}(e(t))$ denotes the approximation error of the ANN. The parameter $l$ corresponds to the number of neurons in the hidden layer. The value function approximation is carried out over a compact domain, and the neural network activation functions are assumed to be bounded on this set. Moreover, the approximation residual $\tilde{\varepsilon}(e(t))$ is assumed to be bounded, which is standard in neural-network-based approximate dynamic programming. The use of a single‑hidden‑layer critic neural network is justified by the universal approximation property, which ensures that the optimal cost function can be approximated with arbitrary accuracy on the compact set $\Omega$ given a sufficient number of neurons. Accordingly, the gradient of $J^{*}(e(t))$, denoted as $\triangledown_{e}J^{*}$, is expressed as:
$\triangledown_{e}J^{*}=\triangledown_{e}\varphi^{\top}W^{*}+\triangledown_{e}\tilde{\epsilon}$, (13)
@@
Since the error $\tilde{\epsilon}(e(t))$ is not directly measurable, we aim to estimate the weight vector $W$. Hence, the approximations for $J(e(t))$ and its gradient are rewritten as:
$\hat{J}(e(t))=\varphi(e(t))^{\top}\hat{W}$, (14)
$\triangledown_{e}\hat{J}=\triangledown_{e}\varphi^{\top}\hat{W}$, (15)
@@
$J^{*}-\hat{J}=\varphi^{\top}(t)(W^{*}-\hat{W})+\tilde{\epsilon}(e(t)),$ (16)
@@
$J^{*}=\hat{J}+\tilde{\delta}+\tilde{\epsilon}(e(t)),$ (17)
@@
$\min\hskip 2.84544ptE=\frac{1}{2}\hat{H}^{\top}\hat{H}$, (20)
Here, the objective function presented in (20) is designed to minimize the quadratic Hamiltonian. By minimizing the objective function $E$, we achieve optimized system performance, characterized by reduced control input $\hat{u}$ while ensuring that the weights $\hat{W}$ converge to their optimal values. To counteract potential distributional shifts due to prioritized sampling, importance weighting and dynamic cluster adjustment are integrated into the gradient update process, ensuring unbiased policy optimization.
@@
The Fast Buffer is a small, adaptive component that prioritizes recent experiences aligned with the current policy $\pi_{\text{curr}}$. New system interactions generate samples, denoted as $S_{\text{new}}=(\mathbf{x}_{t},\mathbf{u}_{t})$, which are immediately stored in the Fast Buffer. By focusing on data that reflects recent agent-environment interactions, the Fast Buffer provides low-bias, high-variance samples, allowing rapid adaptation to evolving dynamics and aligning the agent’s behavior with current environmental conditions. Transitions resulting from CBF-modified control actions are identified as safety‑critical experiences. Such transitions typically occur near the boundary of the safe set and are therefore preferentially stored in the Slow‑Buffer. Through the self‑organizing adaptive clustering mechanism, these safety‑relevant samples are preserved and replayed more frequently, enabling the policy to better learn safe behaviors in critical regions of the state space.
@@
where $m_{j}$ denotes the centroid, and $\sigma_{j}^{2}$ represents the variance of cluster $C_{j}$. This metric enables efficient clustering by ensuring each sample’s integration with the most representative cluster, preventing over-segmentation and enhancing the coherence of clustered data points. The formation of clusters is guided by the computed membership strengths. For any sample with a low similarity to all existing clusters (i.e., $\max(\mu_{C}(S_{\text{old}}))\leq\Gamma_{th}$), a new cluster $C_{\text{new}}$ is created in the Cluster Buffer. This new cluster initializes with $m_{\text{new}}=S_{\text{old}}$ and a variance $\sigma_{\text{new}}=\sigma_{0}=0.02$. Conversely, if a sample fits well within an existing cluster $C_{j}$, the cluster center $m_{j}$ is updated to incorporate the new data point:
$m_{j}=\frac{\mathcal{N}_{j}m_{j}+S_{\text{old}}}{\mathcal{N}_{j}+1},$ (22)
@@
Here, $\beta>0$ denotes the variance amplification coefficient, which controls the adaptability of each cluster by increasing its variance when new samples are absorbed. Larger values of $\beta$ allow clusters to adapt more rapidly to changes in the data distribution, while smaller values enforce more conservative updates and preserve cluster compactness.
@@
$\|m_{i}-m_{j}\|<\gamma\max(\sigma_{i},\sigma_{j}),$ (25)
The parameter $\gamma$ is a global and fixed quantity throughout the clustering process. It is derived from a predefined overlap threshold between Gaussian membership functions and therefore provides a universal criterion for determining when two clusters are sufficiently overlapping to be merged. Now if (25) is satisfied, then overlap between dusters $i$ and $j$ is accured and without loosing the generality, let’s assume $\sigma_{i}\geqslant\sigma_{j}$ then the cluster $C_{j}$ will be pruned and they are merged to form a consolidated cluster $C_{\text{merged}}$ with its center recalculated as:
$m_{\text{merged}}=\frac{\mathcal{N}_{i}m_{i}+\mathcal{N}_{j}m_{j}}{\mathcal{N}_{i}+\mathcal{N}_{j}}.$ (26)
@@
Remark: The value of the parameter $\gamma$, which quantifies the overlap between two clusters, is derived as follows. The overlap is determined by the membership strength $\mu_{C_{i}}(m_{j})$ between the centroids $m_{i}$ and $m_{j}$ based on (21). To ensure a sufficient degree of overlap, we impose the following condition,
$\text{max}(\mu_{C_{i}}(m_{i}),\mu_{C_{j}}(m_{j})))\geq\bar{\gamma}=0.95.$
@@
$\frac{\|m_{i}-m_{j}\|^{2}}{2\sigma_{i}^{2}}\geq\ln(\bar{\gamma})\quad\Rightarrow\quad\frac{\|m_{i}-m_{j}\|^{2}}{\sigma_{i}^{2}}\leq-2\ln(\bar{\gamma}).$
Thus, the cluster $C_{j}$ is covered completely with the cluster $C_{i}$ if:
$\|m_{i}-m_{j}\|\leq\sqrt{-2\ln(\bar{\gamma})}\cdot\sigma_{i}.$
@@
Bias Compensation in Non-Uniform Sampling
Non-uniform sampling in SODACER may shift the input distribution of the neural network, potentially introducing bias. To mitigate this, we incorporate importance sampling weights for each sample drawn from the Slow-Buffer and Fast-Buffer, scaled inversely to its sampling probability. Additionally, the self-organizing clustering dynamically adjusts sampling priorities to align with recent experience distributions. During gradient updates with the Sophia optimizer, these weights are applied to correct the expected gradient direction, ensuring unbiased learning despite non-uniform sampling.
@@
The hyperparameters used in the SODACER–Sophia framework are summarized in Table 1. These parameters govern the dual-buffer architecture, clustering mechanism, and optimization process.
Table 1: Hyperparameters of the SODACER–Sophia framework
Parameter Symbol Value Description
Buffer Architecture
Fast-Buffer capacity $N_{\text{fast}}$ 5,0005{,}000 Maximum number of recent experiences
Max clusters in Slow-Buffer $K_{\max}$ $200$ Maximum number of adaptive clusters
Transfer interval $T_{\text{transfer}}$ $100$ steps Period for moving samples to Slow-Buffer
Membership threshold $\Gamma_{\text{th}}$ $0.7$ Minimum membership for cluster assignment
Clustering Parameters
Initial cluster variance $\sigma_{0}$ $0.02$ Initial variance for new clusters
Variance amplification factor $\beta$ $0.1$ Factor for expanding cluster variance
Variance reduction scaling $\rho$ $1.2$ Scaling factor for variance reduction
Cluster merging threshold $\gamma$ $0.32$ Overlap threshold for merging clusters
Narrow cluster threshold $\sigma_{\text{th}}$ $0.01$ Minimum variance to retain a cluster
Optimization Parameters (Sophia)
Learning rate $\eta$ $1\times 10^{-3}$ Step size for weight updates
First moment decay $\beta_{1}$ $0.9$ Decay rate for gradient moving average
Second moment decay $\beta_{2}$ $0.999$ Decay rate for squared gradient average
Numerical stability term $\epsilon_{0}$ $1\times 10^{-8}$ Small constant to avoid division by zero
Discount factor $\nu$ $0.99$ Discount factor in cost function
Convergence Criteria
Weight convergence threshold $\delta$ $1\times 10^{-4}$ Tolerance for weight updates